find the EXACT value of Cos 5pi/12. no calculators allowed. Give answer in fractional form.
Answer by lwsshak3(11628)

find the EXACT value of Cos 5pi/12. no calculators allowed. Give answer in fractional form.
**
cos(5π/12)=cos(2π/12+3π/12)=cos(π/6+π/4)
cos(s+t)=cos s cos t-sin s sin t
cos(π/6+π/4)=cos π/6 cos π/4-sin π/6 sin π/4
=√3/2*√2/2-1/2*√2/2
=√6/4-√2/4=(√6-√2)/4
ans:
Cos 5π/12=(√6-√2)/4